# DecompGAIL: Learning Realistic Traffic Behaviors with Decomposed Multi-Agent Generative Adversarial Imitation Learning

### v1

@@ 1 Introduction
Realistic traffic simulation is a cornerstone for intelligent transportation systems, autonomous driving, and urban mobility planning. High-fidelity simulators enable safe, controlled evaluation of driving policies, stress-testing of safety-critical scenarios, and principled analysis of interventions Wilkie et al. (2015); Li et al. (2017); Chao et al. (2020). Yet achieving realism remains challenging due to the diversity and complexity of human driving behaviors. Prior rule-based models Treiber et al. (2000), while interpretable and efficient, struggle to capture the complexity and diversity of real-world driving. Behavior cloning (BC) Michie et al. (1990)—as used in Bergamini et al. (2021a); Xu et al. (2023); Philion et al. (2023)—casts imitation as supervised learning from expert trajectories, but suffers from covariate shift Ross et al. (2011): the state distribution induced by a learned policy drifts from that of the expert, compounding errors over time.
These limitations have motivated inverse reinforcement learning (IRL) Ziebart et al. (2008), which learns a reward function that explains expert behavior and then trains a policy via reinforcement learning (RL). By aligning expert and learner trajectory distribution Ghasemipour et al. (2019), IRL theoretically addresses covariate shift. Among IRL approaches, GAIL Ho & Ermon (2016) has gained popularity for traffic modeling Bhattacharyya et al. (2018); Behbahani et al. (2019); Bhattacharyya et al. (2019); Zheng et al. (2020); Wei et al. (2021); Koeberle et al. (2022); Xin et al. (2024), as it bypasses explicit reward inference by training a discriminator to provide a surrogate reward that encourages policies to match the expert’s trajectory distribution. However, in multi-agent settings GAIL often suffers from training instability. To mitigate this, prior work Bhattacharyya et al. (2018) introduces parameter sharing for the policy and discriminator, and curriculum learning strategies that gradually increase the number of vehicles Song et al. (2018) or the rollout horizon Behbahani et al. (2019). While these techniques provide partial improvements, they do not target the root cause and therefore fail to fully resolve the instability.
@@ 2 Related Work
Traffic modeling can be naturally framed as a multi-agent imitation learning (IL) task approached via BC or IRL. BC learns an action distribution conditioned on expert states, parameterized deterministically Bergamini et al. (2021b); Kamenev et al. (2022); Guo et al. (2024), as a Gaussian Yan et al. (2023); Guo et al. (2024); Zhou et al. (2024); Lin et al. (2025), or as a categorical distribution Philion et al. (2024); Wu et al. (2024); Hu et al. (2024); Zhao et al. (2024). However, BC suffers from covariate shift. A direct remedy is data augmentation. Offline approaches perturb demonstrations by injecting noise or dropping trajectory segments Philion et al. (2024); Wu et al. (2024); Hu et al. (2024). More recent online approaches, such as CAT-K Zhang et al. (2024) and uniMM Lin et al. (2025), generate augmented trajectories during learning. CAT-K always selects one of the top-$K$ most likely policy actions that moves the agent closest to the expert next state, while uniMM extends this idea to Gaussian mixture policies by choosing the anchor nearest to the expert future trajectory. Nevertheless, these approaches still suffer from distribution shift, since augmentation remains anchored to expert data rather than policy rollouts.
IRL provides a theoretical solution by letting the agent interact with the environment during training. Among IRL methods, GAIL Ho & Ermon (2016) is most widely used in traffic modeling, but performance degrades in multi-agent domains because of the nonstationary training environment. To stabilize training, prior works Bhattacharyya et al. (2018); Behbahani et al. (2019); Bhattacharyya et al. (2019); Zheng et al. (2020); Wei et al. (2021); Koeberle et al. (2022); Xin et al. (2024) often employ parameter sharing, where all agents use the same policy and discriminator network, thereby pooling experience across agents. Curriculum learning is also applied, gradually increasing the number of vehicles Song et al. (2018) or the rollout horizon Behbahani et al. (2019). However, these methods still produce undesirable behaviors such as off-map driving, collisions, and abrupt braking Bhattacharyya et al. (2019). To mitigate this, some works add hand-crafted rewards alongside adversarial rewards Bhattacharyya et al. (2019); Xin et al. (2024), or purely engineered objectives Peng et al. (2024); Chen et al. (2025). Yet such hand-crafted rewards introduce prior bias and may fail to capture the complexity of human driving behavior.
@@
Beyond traffic modeling, multi-agent GAIL has been explored in other domains. However, most approaches Song et al. (2018); Yu et al. (2019); Zhang et al. (2021) are evaluated in small-scale settings with few agents, where spurious cross-agent interactions are weak. Multi-agent GAIL Song et al. (2018) employs a centralized discriminator that produces a single shared reward, complicating credit assignment and hindering scalability. MA-DAAC Jeon et al. (2020) combines decentralized actor–critic training with adversarial learning, improving stability in cooperative tasks but still struggling to disentangle individual contributions. BM3IL Yang et al. (2020) enhances scalability via a mean-field approximation that averages other agents’ actions as discriminator input, which reduces variance but limits fine-grained interaction modeling. Other works add regularization (e.g., Wasserstein objectives Arjovsky et al. (2017)) to curb discriminator overfitting, at the cost of expressiveness. To our knowledge, prior work has not directly analyzed the root cause of instability in multi-agent GAIL.
@@ 3 Background
A Markov game Littman (1994) is an extension of the Markov Decision Process (MDP) to multiple agents. A Markov game with $N$ agents is defined by the tuple $\langle\mathcal{S},\{\mathcal{A}_{i}\}_{i=1}^{N},P,\{r_{i}\}_{i=1}^{N},\gamma\rangle$, where $\mathcal{S}$ is the state space, $\mathcal{A}_{i}$ is the action space of agent $i$, $P:\mathcal{S}\times\mathcal{A}_{1}\times\dots\times\mathcal{A}_{N}\rightarrow\Delta(\mathcal{S})$ is the transition function, $r_{i}:\mathcal{S}\times\mathcal{A}_{1}\times\dots\times\mathcal{A}_{N}\rightarrow\mathbb{R}$ is the reward function for agent $i$, and $\gamma\in[0,1)$ is the discount factor. Each agent $i$ is equipped with a policy $\pi_{i}:\mathcal{S}\rightarrow\Delta(\mathcal{A}_{i})$, which maps states to distributions over actions. For notational convenience, we use bold symbols to denote joint quantities; for example, $\bm{\pi}$ denotes the joint policy, $\bm{a}$} the joint action.
@@ 4.1 Policy Pretraining
We instantiate our framework within SMART Wu et al. (2024), a widely used learning-based traffic model adopted by prior work Zhang et al. (2024; 2025); Ahmadi & Schofield (2025). SMART parameterizes the joint action as independent categorical distributions over a shared vocabulary of motion tokens. At each time step, it predicts every agent’s next-step token distribution conditioned on encoded map tokens $m$ and all agents’ past motion tokens ${\bm{a}}_{<t}:=({\bm{a}}_{0},\dots,{\bm{a}}_{t-1})$:
@@
Specifically, SMART employs a factorized Transformer Ngiam et al. (2021) with multi-head self-attention (MHSA) and multi-head cross-attention (MHCA) to encode map–motion interactions. The map encoder first produces encoded map tokens $m$ by attending over spatial neighboring map tokens with MHSA. Then, for each agent $i$ at time $t$, the motion encoder takes the current motion-token embedding $e_{t}^{i}$ as the query and sequentially applies temporal, map–agent, and agent–agent attention layers:
@@
where RPE denotes relative positional encoding Cui et al. (2022), and ${\mathcal{N}}_{i}$ is the neighborhood set of map or agent tokens determined by a distance threshold. After $L$ stacked layers of temporal, map–agent, and agent–agent attentions, an MLP head outputs the next-token distribution for each agent.
@@
This pretraining step enables the model to imitate expert behavior on observed trajectories. However, during testing rollouts, small prediction errors can accumulate, leading agents into states not encountered in the demonstrations Ross et al. (2011). This distributional shift often results in implausible behaviors such as collisions or off-road driving Tian & Goel (2025).
@@ 4.2 Decomposed GAIL Fine-tuning
To address the covariate shift issue, we apply GAIL to online fine-tune the BC-pretrained policy. A widely adopted practice in prior works Bhattacharyya et al. (2018); Behbahani et al. (2019); Bhattacharyya et al. (2019) is parameter-sharing GAIL (PS-GAIL), which employs a single discriminator shared across all agents. At each time step, the discriminator evaluates every agent’s local observation, and the resulting score is used as the reward for updating the shared policy $\pi_{\theta}$ via single-agent RL. To implement this, one can replace the policy’s action head with a binary classification head; the discriminator score for ego agent $i$ at time $t$ is:
@@
where ${\bm{a}}_{\leq t}^{{\mathcal{N}}_{i}}:=\{{\bm{a}}_{\leq t}^{j}\mid j\in{\mathcal{N}}_{i}\}$ denotes the motion token histories of all $i$’s neighbors.
Despite its simplicity, PS-GAIL is often unstable Bhattacharyya et al. (2018), degrading simulation quality (see fig.˜3). To examine the source, we conceptually decompose the discriminator’s signal into four terms:
@@
Decomposed Discriminator.
where $\cdot$] denotes concatenation. The discriminator loss is
Social Policy Learning.
A straightforward way to optimize a decentralized multi-agent policy is independent PPO (IPPO) de Witt et al. (2020), which treats each agent as if in a single-agent environment. However, optimizing purely individual objectives can degrade the overall realism of the population Schwarting et al. (2019). We therefore define a social reward for agent $i$:
where $r_{t}^{{\mathcal{N}}_{i}}$ aggregates neighborhood rewards and $\lambda_{ij}$ is a distance-decayed weight (analogous to $w_{ij}$). Training then follows IPPO, but with individual rewards replaced by social rewards. Advantages are estimated with Generalized Advantage Estimation (GAE), and each agent’s value function is given by an MLP applied to the policy network’s final agent–agent attention features. To reduce GPU memory, we freeze the map encoder during fine-tuning and share it between the policy and the discriminator. The full procedure is summarized in algorithm˜1. More details about training can be found in appendix˜A.
@@
   Update $\pi_{\theta},V_{\psi}$ using PPO with clipped surrogate and value loss;
@@ 5 Experiments
  • Q2: How does DecompGAIL compare with state-of-the-art baselines in terms of simulation realism?
  • Q3: What is the contribution of each component within DecompGAIL to overall performance?
@@ 5.1 Experimental Setups
We conduct experiments on the Waymo Open Motion Dataset (WOMD) Ettinger et al. (2021), a large-scale benchmark for traffic simulation containing 487k training, 44k validation, and 44k test scenes. Each scenario is recorded at 10 Hz, with 1 second of historical context and 8 seconds of future trajectories for up to 128 traffic participants (vehicles, cyclists, pedestrians), along with high-definition maps. Following CAT-K Zhang et al. (2024), we evaluate on 2% of the validation split (880 of 44,097 scenarios).
@@
We adopt the official metrics from the WOMD Sim Agents 2025 Challenge (WOSAC) Montali et al. (2024):
@@ 5.3 Performance (Q2)
We compare with six tokenized models Pei et al. (2025); Ahmadi & Schofield (2025); Zhang et al. (2025; 2024); Wu et al. (2024); Peng et al. (2025) and a continuous model Lin et al. (2025), spanning: (i) naive BC methods Wu et al. (2024); Zhang et al. (2025); Peng et al. (2025); (ii) BC with data augmentation Zhang et al. (2024); Lin et al. (2025); and (iii) RL-based approaches Pei et al. (2025); Ahmadi & Schofield (2025). Specifically, SMART-R1 Pei et al. (2025) applies R1-style reinforcement fine-tuning. SMART-tiny-RLFTSim Ahmadi & Schofield (2025) directly optimizes the realism meta-metric as a reward for RL fine-tuning the SMART-tiny model Wu et al. (2024). TrajTok Zhang et al. (2025) enhances SMART by introducing a trajectory tokenizer with better coverage, symmetry, and robustness, along with spatial-aware label smoothing. SMART-tiny-CLSFT Zhang et al. (2024) augments training via expert-guided online trajectory generation to mitigate covariate shift. UniMM Lin et al. (2025) employs a continuous Gaussian mixture model with expert-guided data augmentation. InfGen Peng et al. (2025) models the entire scene as a sequence of tokens—including agent states, motion vectors, and traffic signals—and uses a Transformer to autoregressively simulate traffic.

### v2

@@ 1 Introduction
Realistic traffic simulation is a cornerstone for intelligent transportation systems, autonomous driving, and urban mobility planning. High-fidelity simulators enable safe, controlled evaluation of driving policies, stress-testing of safety-critical scenarios, and principled analysis of interventions (Wilkie et al., 2015; Li et al., 2017; Pan et al., 2025). Yet achieving realism remains challenging due to the diversity and complexity of human driving behaviors. Prior rule-based models (Treiber et al., 2000; Zhang et al., 2021b), while interpretable and efficient, struggle to capture the complexity and diversity of real-world driving. Behavior cloning (BC) (Michie et al., 1990)—as used in Bergamini et al. (2021); Xu et al. (2023); Philion et al. (2024)—casts imitation as supervised learning from expert trajectories, but suffers from covariate shift (Ross et al., 2011): the state distribution induced by a learned policy drifts from that of the expert, compounding errors over time.
These limitations have motivated inverse reinforcement learning (IRL) (Ziebart et al., 2008), which learns a reward function that explains expert behavior and then trains a policy via reinforcement learning (RL). By aligning expert and learner trajectory distribution (Ghasemipour et al., 2019), IRL theoretically addresses covariate shift. Among IRL approaches, GAIL (Ho and Ermon, 2016) has gained popularity for traffic modeling (Bhattacharyya et al., 2018; Behbahani et al., 2019; Bhattacharyya et al., 2019; Zheng et al., 2020; Wei et al., 2021; Koeberle et al., 2022; Xin et al., 2024), as it bypasses explicit reward inference by training a discriminator to provide a surrogate reward that encourages policies to match the expert’s trajectory distribution. However, in multi-agent settings GAIL often suffers from training instability. To mitigate this, prior work (Bhattacharyya et al., 2018) introduces parameter sharing for the policy and discriminator, and curriculum learning strategies that gradually increase the number of vehicles (Song et al., 2018) or the rollout horizon (Behbahani et al., 2019). While these techniques provide partial improvements, they do not target the root cause and therefore fail to fully resolve the instability.
@@ 2 Related Work
Traffic modeling can be naturally framed as a multi-agent imitation learning (IL) task (Brown et al., 2020), approached via BC or IRL. BC learns an action distribution conditioned on expert states, parameterized deterministically (Bergamini et al., 2021; Kamenev et al., 2022), as a Gaussian (Yan et al., 2023; Guo et al., 2024; Zhou et al., 2024; Lin et al., 2025), or as a categorical distribution (Philion et al., 2024; Wu et al., 2024; Hu et al., 2024; Zhao et al., 2024). However, BC suffers from covariate shift Guo et al. (2023). A direct remedy is data augmentation. Offline approaches perturb demonstrations by injecting noise or dropping trajectory segments (Philion et al., 2024; Wu et al., 2024; Hu et al., 2024). More recent online approaches, such as CAT-K (Zhang et al., 2025a) and uniMM (Lin et al., 2025), generate augmented trajectories during learning. CAT-K always selects one of the top-$K$ most likely policy actions that moves the agent closest to the expert next state, while uniMM extends this idea to Gaussian mixture policies by choosing the anchor nearest to the expert future trajectory. Nevertheless, these approaches still suffer from distribution shift, since augmentation remains anchored to expert data rather than policy rollouts.
IRL provides a theoretical solution by letting the agent interact with the environment during training. Among IRL methods, GAIL (Ho and Ermon, 2016) is most widely used in driving behavior modeling (Kuefler et al., 2017), but performance degrades in multi-agent domains because of the nonstationary training environment. To stabilize training, prior works (Bhattacharyya et al., 2018; Behbahani et al., 2019; Bhattacharyya et al., 2019; Zheng et al., 2020; Wei et al., 2021; Koeberle et al., 2022; Xin et al., 2024) often employ parameter sharing, where all agents use the same policy and discriminator network, thereby pooling experience across agents. Curriculum learning is also applied, gradually increasing the number of vehicles (Song et al., 2018; Chen et al., 2022) or the rollout horizon (Behbahani et al., 2019). However, these methods still produce undesirable behaviors such as off-map driving, collisions, and abrupt braking (Bhattacharyya et al., 2019). To mitigate this, some works add hand-crafted rewards alongside adversarial rewards (Bhattacharyya et al., 2019; Xin et al., 2024; Bhattacharyya et al., 2022), or purely engineered objectives (Peng et al., 2024; Chen et al., 2025). Yet such hand-crafted rewards introduce prior bias and may fail to capture the complexity of human driving behavior.
@@
Beyond traffic modeling, multi-agent GAIL has been explored in other domains. However, most approaches (Song et al., 2018; Yu et al., 2019; Zhang et al., 2021a) are evaluated in small-scale settings with few agents, where spurious cross-agent interactions are weak. Multi-agent GAIL (Song et al., 2018) employs a centralized discriminator that produces a single shared reward, complicating credit assignment and hindering scalability. MA-DAAC (Jeon et al., 2020) combines decentralized actor–critic training with adversarial learning, improving stability in cooperative tasks but still struggling to disentangle individual contributions. BM3IL (Yang et al., 2020) enhances scalability via a mean-field approximation that averages other agents’ actions as discriminator input, which reduces variance but limits fine-grained interaction modeling. Other works add regularization (e.g., Wasserstein objectives (Arjovsky et al., 2017)) to curb discriminator overfitting, at the cost of expressiveness. To our knowledge, prior work has not directly analyzed the root cause of instability in multi-agent GAIL.
@@ 3 Background
A Markov game (Littman, 1994) is an extension of the Markov Decision Process (MDP) to multiple agents. A Markov game with $N$ agents is defined by the tuple $\langle\mathcal{S},\{\mathcal{A}_{i}\}_{i=1}^{N},P,\{r_{i}\}_{i=1}^{N},\gamma\rangle$, where $\mathcal{S}$ is the state space, $\mathcal{A}_{i}$ is the action space of agent $i$, $P:\mathcal{S}\times\mathcal{A}_{1}\times\dots\times\mathcal{A}_{N}\rightarrow\Delta(\mathcal{S})$ is the transition function, $r_{i}:\mathcal{S}\times\mathcal{A}_{1}\times\dots\times\mathcal{A}_{N}\rightarrow\mathbb{R}$ is the reward function for agent $i$, and $\gamma\in[0,1)$ is the discount factor. Each agent $i$ is equipped with a policy $\pi_{i}:\mathcal{S}\rightarrow\Delta(\mathcal{A}_{i})$, which maps states to distributions over actions. For notational convenience, we use bold symbols to denote joint quantities; for example, $\bm{\pi}$ denotes the joint policy, $\bm{a}$} the joint action.
@@ 4.1 Policy Pretraining
We instantiate our framework within SMART (Wu et al., 2024), a widely used learning-based traffic model adopted by prior work (Zhang et al., 2025a; b; Ahmadi and Schofield, 2025). SMART parameterizes the joint action as independent categorical distributions over a shared vocabulary of motion tokens. At each time step, it predicts every agent’s next-step token distribution conditioned on encoded map tokens $m$ and all agents’ past motion tokens ${\bm{a}}_{<t}:=({\bm{a}}_{0},\dots,{\bm{a}}_{t-1})$:
@@
Specifically, SMART employs a factorized Transformer (Ngiam et al., 2021) with multi-head self-attention (MHSA) and multi-head cross-attention (MHCA) to encode map–motion interactions. The map encoder first produces encoded map tokens $m$ by attending over spatial neighboring map tokens with MHSA. Then, for each agent $i$ at time $t$, the motion encoder takes the current motion-token embedding $e_{t}^{i}$ as the query and sequentially applies temporal, map–agent, and agent–agent attention layers:
@@
where RPE denotes relative positional encoding (Cui et al., 2022), and ${\mathcal{N}}_{i}$ is the neighborhood set of map or agent tokens determined by a distance threshold. After $L$ stacked layers of temporal, map–agent, and agent–agent attentions, an MLP head outputs the next-token distribution for each agent.
@@
This pretraining step enables the model to imitate expert behavior on observed trajectories. However, during testing rollouts, small prediction errors can accumulate, leading agents into states not encountered in the demonstrations (Ross et al., 2011). This distributional shift often results in implausible behaviors such as collisions or off-road driving (Tian and Goel, 2025).
@@ 4.2 Decomposed GAIL Fine-tuning
To address the covariate shift issue, we apply GAIL to online fine-tune the BC-pretrained policy. A widely adopted practice in prior works (Bhattacharyya et al., 2018; Behbahani et al., 2019; Bhattacharyya et al., 2019) is parameter-sharing GAIL (PS-GAIL), which employs a single discriminator shared across all agents. At each time step, the discriminator evaluates every agent’s local observation, and the resulting score is used as the reward for updating the shared policy $\pi_{\theta}$ via single-agent RL. To implement this, one can replace the policy’s action head with a binary classification head; the discriminator score for each agent $i$ at time $t$ is:
@@
where ${\bm{a}}_{\leq t}^{{\mathcal{N}}_{i}}:=\{{\bm{a}}_{\leq t}^{j}\mid j\in{\mathcal{N}}_{i}\}$ denotes the motion token histories of all $i$’s neighbors. When computing the reward for each agent $i$, we treat that agent as the ego agent.
Despite its simplicity, PS-GAIL is often unstable (Bhattacharyya et al., 2018), degrading simulation quality (see fig.˜3). To examine the source, we conceptually decompose the discriminator’s signal into four terms:
@@
   Update $\pi_{\theta},V_{\psi}$ using PPO with clipped surrogate and value loss and BC loss;
@@
Decomposed Discriminator.
where $\cdot$] denotes concatenation. The discriminator loss is
Social Policy Learning.
A straightforward way to optimize a decentralized multi-agent policy is independent PPO (IPPO) (de Witt et al., 2020), which treats each agent as if in a single-agent environment. However, optimizing purely individual objectives can degrade the overall realism of the population (Schwarting et al., 2019). We therefore define a social reward for agent $i$:
where $r_{t}^{{\mathcal{N}}_{i}}$ aggregates neighborhood rewards and $\lambda_{ij}$ is a distance-decayed weight (analogous to $w_{ij}$). Training then follows IPPO, but with individual rewards replaced by social rewards. Advantages are estimated with Generalized Advantage Estimation (GAE), and each agent’s value function is given by an MLP applied to the policy network’s final agent–agent attention features. To improve stability, we combine the PPO loss and BC loss. The full procedure is summarized in algorithm˜1. More details about training can be found in appendix˜A.
@@ 5 Experiments
  • Q2: How does DecompGAIL compare with state-of-the-art baselines in terms of realism?
  • Q3: What is the contribution of each component to overall performance?
@@ 5.1 Experimental Setups
We conduct experiments on the Waymo Open Motion Dataset (WOMD) (Ettinger et al., 2021), a large-scale benchmark for traffic simulation containing 487k training, 44k validation, and 44k test scenes. Each scenario is recorded at 10 Hz, with 1 second of historical context and 8 seconds of future trajectories for up to 128 traffic participants (vehicles, cyclists, pedestrians), along with high-definition maps. Following CAT-K (Zhang et al., 2025a), we evaluate on 2% of the validation split (880 of 44,097 scenarios).
@@
We adopt the official metrics from the WOMD Sim Agents 2025 Challenge (WOSAC) (Montali et al., 2024):
@@ 5.2 Stability (Q1)
Table 1: Results on the WOSAC 2025 leaderboard test split. (Waymo, 2025)
SMART-tiny-DecompGAIL (ours) 0.7864 0.4919 0.8152 0.9176 1.4209
SMART-R1 (Pei et al., 2025) 0.7858 0.4944 0.8110 0.9201 1.2885
SMART-tiny-RLFTSim (Ahmadi and Schofield, 2025) 0.7857 0.4927 0.8129 0.9183 1.3252
TrajTok (Zhang et al., 2025b) 0.7852 0.4887 0.8116 0.9207 1.3179
SMART-tiny-CLSFT (Zhang et al., 2025a) 0.7846 0.4931 0.8106 0.9177 1.3065
UniMM (Lin et al., 2025) 0.7829 0.4914 0.8089 0.9161 1.2949
SMART-tiny (Wu et al., 2024) 0.7814 0.4854 0.8089 0.9153 1.3931
LLM2AD (Wang et al., 2025) 0.7779 0.4846 0.8048 0.9109 1.2827
InfGen (Peng et al., 2025) 0.7731 0.4493 0.8084 0.9127 1.4252
@@ 5.3 Performance (Q2)
We compare both tokenized models (Pei et al., 2025; Ahmadi and Schofield, 2025; Zhang et al., 2025b; a; Wu et al., 2024; Peng et al., 2025; Wang et al., 2025) and continuous model (Lin et al., 2025), covering a broad spectrum of learning paradigms:
  • Naive BC (Wu et al., 2024; Zhang et al., 2025b; Peng et al., 2025);
  • BC with data augmentation (Zhang et al., 2025a; Lin et al., 2025);
  • RL–based finetuning approaches (Wang et al., 2025; Pei et al., 2025; Ahmadi and Schofield, 2025).
Specifically, SMART-R1 (Pei et al., 2025) applies R1-style reinforcement fine-tuning. SMART-tiny-RLFTSim (Ahmadi and Schofield, 2025) directly optimizes the realism meta-metric as a reward for RL fine-tuning the SMART-tiny model (Wu et al., 2024). TrajTok (Zhang et al., 2025b) enhances SMART by introducing a trajectory tokenizer with better coverage, symmetry, and robustness, along with spatial-aware label smoothing. SMART-tiny-CLSFT (Zhang et al., 2025a) augments training via expert-guided online trajectory generation to mitigate covariate shift. UniMM (Lin et al., 2025) employs a continuous Gaussian mixture model with expert-guided data augmentation. LLM2AD (Wang et al., 2025) integrates GRPO post-training with test-time search and clustering on a larger SMART backbone equipped with enhanced tokenization and positional encoding. InfGen (Peng et al., 2025) models the entire scene as a sequence of tokens—including agent states, motion vectors, and traffic signals—and uses a Transformer to autoregressively simulate traffic.